OpenStudy (anonymous):

calculate the number of bytes required to store illumination and express the exact number of bytes in 1 tb as power of 2

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I'm sorry, I don't know your answer

OpenStudy (michele_laino):

please wait I ask to another tutor

OpenStudy (anonymous):

ok. thanks

OpenStudy (michele_laino):

@dan815 can you help here please?

OpenStudy (rational):

1 terabyte = 2^10 gigabyte = 2^10*2^30 byte = 2^40 byte

OpenStudy (anonymous):

thanks @rational but what about ....calculate the number of bytes required to store illumination

OpenStudy (rational):

are you storing in ASCII format ?

OpenStudy (anonymous):

yes

OpenStudy (rational):

count the number of characters in that word

OpenStudy (anonymous):

they are 12

OpenStudy (rational):

how much memory each character takes ?

OpenStudy (anonymous):

1 byte each

OpenStudy (rational):

so how many bytes 12 characters take

OpenStudy (anonymous):

12 bytes

OpenStudy (rational):

simple arithmetic

OpenStudy (anonymous):

waw thanks

OpenStudy (rational):

yw